Question 11
A person standing on level beach throws a ball with a velocity of 22 m/s at an angle of 30o above the horizontal. How far away is the ball when it hits the beach and what is the time of flight?
1
Expert's answer
2019-07-19T14:09:41-0400

"Vx=Vi\times Cos\theta"

This is the horizontal component of velocity

"Vx=22\times Cos30^0"

"Vx=19.05\frac{m}{s}"

Now the vertical component of velocity"Vy=Vi\times Sin\theta"

"Vy=22*Sin30^0"

"Vy=11\frac{m}{s}"

Time of flight

"t=\frac{2Vy}{a}"

t=2*11/9.81

"t=2.24sec"

Distance traveled when hit the beach again

D=Horizontal velocity multiply by time of flight

D=19.05x2.24

D=42.76meter